Write the exponential function y=ab^x whose graph passes through the given points: (1,6), (2,36)
Answer by scott8148(6628)

substituting __ 6=ab

substituting __ 36=ab^2

dividing __ 36/6=(ab^2)/(ab) __ 6=b

substituting __ 6=a(6) __ 1=a